A forensic scientist analyzes a sample for sodium arsenate, Na 3 AsO 4 (s), a source of arsenic. Calculate

the percentage composition of sodium arsenate.


1
Expert's answer
2022-06-06T17:04:09-0400

Mr(Na) = 23 g/mol

Mr(As) = 75 g/mol

Mr(O) = 16 g/mol

Mr(Na3AsO4) = 208 g/mol



"Na= \frac{23*3}{208}*100= 33,17%" %


"As= \frac {75}{208}*100= 36,06%" %


"O= \frac {16*4}{208}*100= 30,77%" %